2x+y=-2 and 5x+3y=-8 solved by using substitution method
Show work

Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

2x+y=-2 and 5x+3y=-8 solved by using substitution method
***
2x+y=-2
y=-2x-2
..
5x+3y=-8
sub y from 1st equation
5x+3(-2x-2)=-8
5x-6x-6=-8
add 6 to both sides and add x-terms
-x=-2
divide both sides by (-1)
x=2
y=-2x-2=-4-2=-6

Answer by stanbon(75887)

2x+y=-2 and 5x+3y=-8 solved by using substitution method
--------
Solve each for "y"::
y = -2x-2
y = (-5x-8)/3
-----
Substitute and solve for "x"::
----
-2x-2 = (-5x-8)/3
----
-6x-6 = -5x-8
x = 2
-----------
Solve for "y"::
y = -2x-2
y = -4-2 = -6
------------------
Cheers,
Stan H.
-----------